one mole of N2O4 (g) at 300k is kept in a closed container under 1 atm. it is heated to 600k when 20% by mass of N2O4 decomposes to NO2(g). the resultant pressure is?

Grade:12

1 Answers

Aman Bansal
11 years ago

Dear Shubham,

ans is (b) 2.4 atm

 

as temp is doubled pressure will also doubled ,as vol is constant

 

at equilibrium pressure due to N2O4 = 2 - 2(0.2) = 1.6 atm

 

pressure due to NO2 = 4*(0.2) = 0.8 atm

 

so total pressure = 1.6 0.8  =  2.4 atm

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert